a radioactive substance decays to 1/32 of its initial value in 25 days calculate its half life

Grade:12

1 Answers

Nicho priyatham
8 years ago
the amount of  substance left after ‘n’ half lives is (1/2)n(intial amount)
N=(1/2)n N 
in the given problem n is 5 i.e (1/2)5=1/32 
so 5 have lives is 25 days
and 1 half life is 25/5=5days
so half life is 5 days
PLEASE APROVE